A winter coat manufacturer is deciding whether to manufacture a puffer coat or a parka. For the puffer coat the cost of producing each item is $50.55 and they will sell each item to a reseller for $167.48. For the parka the cost of producing each item is $65.66 but they will sell each item to a reseller for $200.98.


What is the contribution rate for manufacturing the puffer coats? Your answer should be accurate to at least two decimal places.

What is the contribution rate for manufacturing the parkas? Your answer should be accurate to at least two decimal places.

Which type of coat has the better contribution rate?

Pleas
Answer #1
contribution rate for manufacturing the puffer coats = ( Selling price - Cost per unit ) / Selling price = ( 167.48 - 50.55 ) / 167.48 = 69.82%
contribution rate for manufacturing the parkas = ( Selling price - Cost per unit ) / Selling price = ( 200.98 - 65.66 ) / 200.98 = 67.33%
Which type of coat has the better contribution rate?
Answer : puffer coats [ as the contribution rate is higher ]